The IFO is ready for 3F DRMI comissioning
Pump down reached "vacuum normal" state. IFO _P1 pressure 1e-4 torr in 8 hrs actual pumping time
PSL shutter is opened.
We stopped pumping just short of 3 hours at 320 Torr. Pumping speed was 2.7 Torr / min with partially closed RV1 and butterfly valve/
RP1&3 roughing pump hose is disconnected. Butterfly valve removed. The vac envelope is closed.
This is our second stop. I will be back this afternoon. IFO P1 3.5 Torr
I've installed a new 2pin lemo cable going from the CM servo out to in2 of the MC servo board, and removed the temporary BNC. I used some electrical tape to give the cable some thickness where the lemo head screws on to try to strain relieve the solder joints; hopefully this cable is more robust than the last.
I put an excitation into the CM board, and saw it come out of MC_F, so I think we're set.
Q checked the earth quake stops of SRM and we put the ITMY & BS doors on.
Photos have been taken of the ITMY chamber, and uploaded to picasa. Here's a slideshow:
I closed the PSL shutter as we didn't want to burn the mirror surface when we are not working.
Everything is aligned, AS and POY make it out of vacuum unclipped, OSEM readings look good.
I set up the SRM oplev, centered all oplevs.
Tomorrow, we just have to take pictures of the ITMY chamber before we put the heavy doors on.
We have completed the above points; the ITMY table is still level.
Despite what the wiki says, the SRM LR OSEM open voltage is ~1.97V instead of ~1.64, so we shot for half of that.
The in-air steering of the SRM oplev return beam needs adjustment. I'll estimate the beam path length when I'm taking pictures and closing up.
Left to do:
Related In-Air work:
The RGA time stamp was correct last at 20140527
Rga stopped scanning at 20140530
POY has >2 inches of clearance from the SRM cage.
Distance reconstruction indicates an SRC length of 5399mm, which was exactly our target.
Today so far:
Here's my quick brain dump of things to do before we can pump down (anyone see anything missing?):
I figured out that didn't change the initial guess for the fit routine in Gabriele's code. I also changed the fminsearch criteria to least squares fitting, instead of minimax. The consistency checks now look just as good as the previous time we did these kind of measurements, no disagreements bigger than 1.6mm.
Thus, the current estimate of the SRC length after yesterday's motions is 5402mm, where we desire 5399mm. So, we will try to move SRM 3mm closer to SR2, after confirming that we are not clipping the POY beam. After all that, we will level the table.
Larry and Nicolas
Larry's transfer function measurements suddenly started returning 0dB 0degrees when before there was some fake filter in the C1:ALS-OFFSETTER2 filter bank.
We looked in the filter bank and his filter was gone. So I created a new filter called LARRYP in FM2. We also disabled the output so he could drive the filter bank and test his TF code.
ITMY in vac table needs leveling.
[q, Jenne, Manasa]
I think I found out why rossa was mad.
An apt-get update on the 18th downloaded kernel 2.6.32-65-generic, so 2.6.32-58-generic, which what was previously chosen as a working kernel, had moved down in the grub ordering.
It turns out the grub configuration accepts strings, so I changed it to GRUB_DEFAULT="Ubuntu, with Linux 2.6.32-58-generic", ran sudo update-grub, and Rossa now seems to boot happily.
GRUB_DEFAULT="Ubuntu, with Linux 2.6.32-58-generic",
[Zach, Jenne, Steve]
This work happened on Tuesday. Bad Jenne for forgetting to elog it!
Zach brought the 40m's seismometers back (one Guralp and one T-240). We have set the seismometers on their slabs. Also, we ran the T240 cable from 1X5 over to the vertex slab. Also, also, Zach and Steve mounted the T-240 readout box in the 1X5 rack. We have not yet hooked it up to power, although there are fused power blocks available on that rack.
So, the T-240 box needs power, and then we need to connect the seismometers to their respective boxes. Also, we need to run medium-short BNC cables from the T-240 readout box to the PEM AA board over in 1X7.
I looked at the CAD layout and it seems like we will clearly be clipping POY if we move SRM by 7.5cm. Since POY is not visible at low power, we cannot be sure about the clipping.
I was bad and forgot to elog this yesterday (bad grad student!), but I setup a laser pointer to show us where the POY beam is.
To do this, I removed the tiny mirror that sends the beam to the POY RF PD (so we do not have POY to lock the Yarm right now. I think Q has successfully been using AS). The laser pointer goes through 2 temporary steering mirrors, then passes through the place that the tiny mirror usually sits, and then travels along the POY path into the vacuum system. The idea here is that we should be able to adjust the laser pointer and the temp steering mirrors, and not touch any of the actual POY mirrors, but still get the green beam to go all the way to ITMY. Yesterday I confirmed that the laser pointer was hitting the in-vac POY pickoff mirror, and today Q and Manasa are doing final adjustment to get the beam all the way to the ITM.
[Steve, EricQ, Jenne]
ITMY and BS heavy doors are off, light doors are on. Q is aligning the IFO.
Ottavia was having some severe interaction latency today. Xorg was taking up >90% of the CPU, just sitting around. The machine was logged in to a desktop session with lots of graphical effects turned on. I changed the system default session to "gnome-fallback" in /etc/lightdm/lightdm.conf, which was already set as the default for controls, but wouldn't get chosen for the autologin that happens on boot.
Hopefully this helps ottavia stay usable...
Interferometer alignment is restored
ASS has been run on each arm, recycling mirrors were aligned by overlapping on AS camera.
ETMY was not getting its ASC pitch and yaw signals. C1SCY had a red RFM bit (although, it still does now...)
I took a look at the c1rfm simulink diagram and found that C1RFM had an RFM block called C1:RFM-TST_ETMY_[PIT/YAW] and C1SCY had one called C1:TST-SCY_ETMY_[PIT/YAW].
It seems that C1TST was illegally being used in a real signal chain, and Jenne's recent work with c1tst broke it. I renamed the channels in C1RFM and C1SCY to C1:RFM-SCY_ETMY_[PIT/YAW], saved, compiled, installed, restarted. All was well.
There are still some in SCY that have this TST stuff going on, however. They have to do with ALS, it seems, but are SHMEM blocks, not RFM. Namely:
The vent is completed. ITMX was kicked up accidentally. Valve configuration: chamber open, RGA is pumped through VM2 Maglev
We should have a plan B before we move everything. I suggest we move a combination of SRM and SR2 to get the desired SRC length.
Moving SR2 will require extra effort to walk the beam unclipped through all the 6 output steering mirrors that follow; but there will be little room for error if we use irides to propagate the beam through the first 4 mirrors that are in the BS and ITMY chamber.
Jam nuts checked. Oplev servos turned off. Particle count checked. Vertex crane functionality checked.
As Q mentioned in elog 10527, (prompted by Koji's email this afternoon) we are prepping the IFO for vent. Here is a copy of the pre-vent checklist from the wiki, updated as we work:
1 & 2: Locked arms on IR, ran ASS. Unlocked IFO, aligned PRM for good POP flashes, aligned SRM for symmetric AS flashes. Aligned all oplevs. Used PZTs to align Xgreen to arm. Used knobs to align Ygreen to arm. With PS:L green shutter closed, Xgreen = 0.520, Ygreen = 0.680.
3: Moved MC servo output cable that goes to ADC from OUT2 (which we had been using for monitoring AO path signals) back to its usual OUT1 (which is MC_L). This is used in the spot position measurement script. Spots at: [2.32, -0.50, 1.97, -1.11, 0.26, -1.86] mm.
4: Done -Q
5: Removed a PD that was monitoring the light coming backwards through the Faraday that sits just after the laser, just in case (confirmed that beam dump behind PD was catching beam). Other port of PBS just had regular black hole dump. Adjusted half wave plate until we had ~90mW just before injection into the vacuum.
6: Completed. Locked MC manually at transmission of ~1150, but low power autolocker isn't working. This isn't a critical thing, and can be fixed at any point during the vent. -Q
7: Shutters closed. Ready for Steve to check nuts and begin venting! -Q
Rather than using a CAD drawing, I used Gabriele's code from ELOG 9590 to try and judge if we could shorten the SRC by the appropriate length, without clipping the SR3-SR2 beam.
Specifically, I used these lines:
% Move SRM 7.5 towards SR2, parallel to beam
dAS = BS2-AS; % Vector from SRM to SR2
dASmag = sqrt(dAS(1)^2+dAS(2)^2);
dMove = delta*dAS/dASmag; % delta times unit vector
CS = CS+dMove;
As a reminder, Gabriele's code used the following logic:
In my opinion, this is the best estimate of beam trajectory that we currently have.
Thus, from looking at the plot above, I claim we can correct the SRC length without clipping the beam by moving the SRM forward by the required 7.5cm.
Although the measured distance may be off on the order of a cm (since our PRC correction had a 0.5cm disagreement between interferometric and hand distance measurements), this will nevertheless markedly improve our 3F DRMI sensing, based on my previous ELOG.
Hence, given our discussions last week, Jenne and I will proceed to ready the interferometer for venting in the morning, by following the vent checklist.
Our sole objective for this vent is this move of the SRM.
Steve, please check the jam nuts, and begin the vent when you get in. Thanks!
The room temp drops 1 degree C on the 4th day. The weather has changed.
TEST QPD sn 222 was calibrated with 1103P directly looking into it from 1 m. ND2 filter was on the qpd.
Discontinuities / glitches could be seen in the CM board fast output when MC board gains were changed, which isn't so nice. Incidentally, I notice now that each lock loss corresponded to a step of AO gain on the CM board.
Back in May I looked at all the glitches that happen when we change the AO gain slider on the CM board - see elog 9938. I wonder if the MC IN2 gain slider has the same issues. I think I'll look at this this afternoon. Maybe we can set the CM board gain someplace, and just use the MC IN2 slider (if it's not as glitchy) for the delicate part where we're just about to cross unity, and then later we can again use the CM board's AO gain.
EDIT: Yes, the glitches on the CM board AO path are *much* bigger, and more frequent. Interestingly, the biggest glitches were every 4 dB. When I went from -29 to -28, again from -25 to -24, -21 to -20, etc. I saw the largest glitches on the MC IN2 slider going -29 to -28 and -17 to -16, but if there were small glitches at other transitions, they didn't hit my trigger levels. I think next time I try engaging the AO path I'll try to do the delicate stuff by upping the MC IN2 gain rather than the CM board AO gain.
More AO efforts. No huge news.
Came at AO from each side. For each sign, I lost lock just a few dB from the AO portion of the loop crossing unity gain. Both attempts were about arm powers of 1, which should correspond to ~300pm CARM offset, which I have simulated the crossover as possible with my current loop models (including latest MC loop). The gain steps were usually 6dB in between measurements.
Positive polarity on CM board screen:
I made it to +5 dB of the last plot here, but the 6th broke it open. Gains on CM In2, CM AO, and MC In2 were -6, -4, -2 on that last, lock breaking, step.
Negative polarity on CM board screen:
Lost it just 2dB above the last trace. Gains were -6, +1, -2 (So, overall 5dB higher than the other polarization)
Many things happened in between these two lock stretches, but I'm not sure what may or may not have affected things. They include:
The location of the CARM resonance peak lines up with my simulation, which is good, but there appears to be less phase than expected... I tried making sure that we don't have any whitening uncompensated for, but it looked ok. All my AO path loop model contains is the CM board TF (measured and fitted), the IMC seen as an actuator(measured and fitted), and the REFLDC optical TF (simulated in MIST). Maybe the DC path of whatever diode this is coming from needs to be included...
We pulled the old 2-pin lemo cable after I had a look at the connectors. When I unscrewed the connector on the MC side, one of the wires came off. I suspect that it was still hanging on a bit, but my torquing it finally killed it.
We pulled the cable with the idea of resoldering the connectors, but there are at least 2 places where the cable has been squished enough that the shielding or the inner wires are exposed. These places aren't near enough the ends to just cut the cable short.
Downs doesn't have a spool of shielded twisted single-pair cable, so Todd is going to get me the part number for the cable they use, and I've asked Steve to order it tomorrow.
For now, we will continue using the BNC cable that we installed last night - I don't think it's worth resoldering and putting in a crappy 2-pin lemo cable that we'll just throw out in a week.
2005 ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW
Sep. 28, 2006 optical lever noise budget with DC readout in 40m, LIGO- T060234-00-R, Reinecke & Rana
May 22, 2007 BS, SRM & PRM He Ne 1103P takes over from diode
May 29, 2007 low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007
Nov 30, 2007 Uniphase 1103P divergence measured
Nov. 30, 2007 ETMX old Uniphase 1103P from 2002 dies: .............., running time not known......~3-5 years?
May 19, 2008 ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~ ?
Oct. 2, 2008 ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier
JDSU 1103P were replaced as follows:
May 11, 2011 ETMX replaced, life time 1,258 days or 3.4 years
May 13, 2014 ETMX , LT 1,098 days or 3 y
May 22, 2012 ETMY, LT 1,464 days or 4 y
Oct. 5, 2011 BS & PRM, LT 4 years, laser in place at 1,037 days or 2.8 y
Sep. 13, 2011 ITMY old 1103P & SRM diode laser replaced by 1125P ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y
June 26, 2013 ITMX 622 days or 1.7 y note: we changed because of beam quality.........................laser in place 420 days or 1.2 y
Sep. 27, 2013 purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )
May 13, 2014 ETMX, .............laser in place 90 d
May 22, 2012 ETMY,
Oct. 7, 2013 ETMY, LT 503 d or 1.4 y............bad beam quality ?
Aug. 8, 2014 ETMY, .............laser in place 425 days or 1.2 y
Sept. 5, 2014 new 1103P, sn P893516 installed at SP table for aLIGO oplev use qualification
Tonight was a night of trying to engage the AO path. The idea was to sit at arm powers of a few on sqrtInvTrans for CARM and ALS for DARM, and try to increase the gain for REFLDC->AO path.
No exciting nit-picky details in locking procedure. Mostly it was just a night of trying many times.
The biggest thing that Q and I found tonight was that the 2-pin lemo cable connecting the CM board's SERVO OUT to the MC board's IN2 is shitty. The symptom that led to this investigation was that I could increase the AO path gain arbitrarily, and have no change in the measured analog CM loop transfer function. We checked that the CM board servo out spit out signals that were roughly what we expected based on our ~2kHz excitation. However, if we look at digitized signals from the MC board, the noise level was very high, with loads of 60Hz lines, and a teensy-tiny signal peak. We put a small drive directly into the MC board and could see that, so we determined that the cable is bad. We have unplugged the white 2-pin lemo, and ran a long BNC cable between the 2 boards. Tomorrow we need to make a new 2-pin lemo cable so that we can have the lower noise differential drive signal.
After putting in the temporary cable, we do see an excitation sent to the CM board showing up after the MC board. For this monitoring, the MC_L cable to the ADC has been borrowed, so instead of being the OUT1, the regular length signal, MC_L is currently the OUT2 monitor right after the board inputs.
At some point in the evening, around 1:15am, ETMX started exhibiting the annoying behavior of wandering off sometimes. I went in and pushed on the SUS cables to the satellite box, and I think it has helped, although I still saw the drift at least once after the cable-squishing.
Other than that, it has just been many trials.
The best was one where I was holding the arm powers around 4, and got the CM board's AO gain to -8 dB and the MC board's IN2 AO gain to -4 dB. I lost lock trying to increase the CM board gain to -7 dB.
I took several transfer functions, and used Q's nifty "SRmeasure" script to gather data, and Q made a plot to see the progress.
TF progress plots:
Time series of that lockloss:
I don't know yet if the polarity of the CM board should be plus or minus. This series was taken with "minus". But, since the phase looked opposite of Q's single arm CM board checkout from several months ago, we did a few trials with the polarity switched to "plus". I thought we weren't getting as high of AO path gains, so I switched back to "minus", but the last few trials didn't get even as far as the plus trials did. So, I still don't know which sign we want.
diag < test.dtt
I have not had any success the past two days in getting an interferometric measurement of the SRC length.
So, the question posed at today's meeting was: "How precisely do we need to change the SRC length to be able to lock the DRMI on 3F"
The two ways I could think to quantify this are:
REFL33 should have its phase set to put PRCL along I, and REFL165 should have SRCL along I, so the simulation result that matters is the angle of MICH in these planes. The cross couplings are then given by the appropriate trigonometric projections. In the following plots, I used 10% as the acceptable cross coupling in either direction.
Code (finesse + pykat + ipython notebook) and plots are attached.
Steve asked about calibrating the QPD, so I set up some new epics records so that we can have calibrated versions of the QPD output.
The new channels are called C1:ASC-TESTQPD_Y_Calc and C1:ASC-TESTQPD_X_Calc for pitch and yaw, respectively.
* I modified /cvs/cds/caltech/target/c1iscaux/QPD.db to add 2 new channels. Since we are currently plugged into the IPPOS channels, I didn't want to modify the units of IPPOS, which is why I created new channels. The new channels are just the IPPOS normalized X and Y channels, multiplied by a calibration factor. Steve has already done a rough calibration for his setup, so I used those numbers (0.15 urad/ct for pitch and 0.25 urad/ct for yaw).
* Rebooted c1iscaux. This required adding it to chiara's /etc/hosts file.
* Added the channels to the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file so that the channels would appear in dataviewer.
* Restarted the framebuilder daqd process.
How to modify the calibration:
1) On a control room workstation, cd /cvs/cds/caltech/target/c1iscaux to get to the right folder. (Note that this is still in the old cvs/cds place, *not* the new opt/rtcds place)
2) open the epics database file by typing sudo emacs QPD.db. Since this is a protected file, you need to use the "sudo" command, and will have to type in the usual controls password.
sudo emacs QPD.db
3) Find the "records" that have the channel names C1:ASC-TESTQPD_Y_Calc and C1:ASC-TESTQPD_X_Calc by scrolling down. (Right now they are on lines #550 and #561 of the text file).
4) For each of these 2 records, modify the calibration in the line that says something like field(CALC,"(A*0.25)"). In this example, the current calibration is 0.25 urad/oldCount. Change the number to the new value.
5) Save the file. If you followed the procedure in step2 and used the emacs program and you can't use the mouse, do the following: Hold down the "ctrl" key. Keeping ctrl pushed down, push the "x" key. Still keeping ctrl pushed down, push the "s" key.
6) Close the file. If you followed the procedure in step2 and used the emacs program and you can't use the mouse, do the following: Hold down the "ctrl" key. Keeping ctrl pushed down, push the "x" key. Still keeping ctrl pushed down, push the "c" key.
7) Reboot the slow computer called c1iscaux. You should be able to do this remotely by typing telnet c1iscaux, and then typing reboot. If that doesn't work, you may have to go into the IFO room and power cycle the crate by turning the key. This computer is in 1Y3, near the bottom.
8) Check that you can see your channels - you should be finished now!
For steps 3 and 4, here is a screenshot of the lines in the text file:
I tried a bunch of times to reduce my CARM offset so I could jump to REFLDC digitally, but I think I'm maybe being a little ambitious with the arm power I'm trying to get to.
I have modified the carm_cm_up script so that it does my new procedure. Everything is the same through locking the PRCL and MICH on 3f. Then it reduces the CARM offset to 1.5 nm. This is where we *used* to transition to sqrtInvTrans. Now I have it going a bit farther to 0.5 nm, and arm powers of about 1 before doing that transition. Also, before it transitions it lowers the CARM gain and engages the 1kHz lowpass in FM9. A gain of about 4 is fine to keep the gain peaking in the CARM loop to only about 10dB, and sets a UGF of 100Hz which is the peak of the phase bubble with the lowpass engaged.
Once I got to this point (several times tonight), I turned on CARM and DARM oscillations and looked at the transfer functions between (CARM and REFLDC) and (DARM and AS55Q). I have 2 DTT templates setup for this, in /users/Templates/PRFPMI. These templates assume that you have your new DARM signal (AS55) going to SRCL_IN1 and your new CARM signal (REFLDC, which is actually REFL11I coming through the CM board) going to MC_IN1.
I'm not sure why I'm losing lock. I don't see anything terribly telling on the time series plots, in particular none of the loops look like they are oscillating. Here is one of the better examples from this evening:
* I realigned the Xgreen on the PSL table (again) to maximize the beatnote amplitude. Y was fine, but X was very poorly overlapped on the camera.
* I put the SR785 back by the LSC rack and plugged it into the CM board for transfer functions. Didn't take any tonight.
* We have a small wishlist for scripting things: (1) DRMI restore script should reset REFL11 to "normal" REFL11. (2) CARM/DARM acquisition restore script should reset REFL11 to REFLDC. (3) CARM/DARM acquisition restore should also set PRMI parameters (as Q noted last week).
Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot
Here's the sequence of the morning so far:
The IFO is still down, as the PMC won't lock without the rack power, and we haven't pinned down the shorting mechanism. We don't want the replacement sorensen to immediately blow when plugged in.
The Uniblitz shutter is insulated at the optical table. I only realized it when it was in my hand. The 3"x2" base plate is black DELRIN !
In the future we must label black Delrin base plate where it is used. Now we have white Delrin and light bran PEEK base plates for the same function.
Our janitor turned off the laser accidentally.
Didn't you take this opportunity to replace the cooling fan of the innolight controller?
The PMC wasn't locking very happily after this. I tweaked the pointing onto the PMC REFL diode, to make sure it was centered, and touched the alignment into the PMC. I also reset the FSS Slow output to zero. It took a little while for the laser to settle in, for some reason, but the transmission is up at 0.80 now.
Tweaked MC2 pointing to get the MC transmission high enough to let WFS kick in, which nicely got the rest of the MC alignment done. After that, I offloaded the WFS into the MC suspensions.
Lastly, I ran the command that Rana posted in ELOG 10391, to set the FSS input offset (From -0.18 to -0.06)